# SPAWNing Structural Priming Predictions from a Cognitively Motivated Parser

### v1

@@ Appendix C Details about the training dataset
Appendix D Details about statistical models
D.1 Model specification
@@
We fit the maximal model by including all by-model-instance and by-item random intercepts and slopes.
D.2 Priors
we fit the model using the following weakly informative prior.
@@ D.2 Priors
This prior assumes that the log odds ratio between priming conditions is most likely to be 0 (i.e. no priming effect) and unlikely to be greater than 4 or less than -4. This assumption is based on a meta-analysis of priming in production studies (mahowald2017) where the log odds ratio between the prime conditions was not greater than 4 in any of the constructions they considered.
D.3 Statistical inferences
@@
From the posteriors of the Bayesian models, we computed 95% credible intervals for P(RRC prime, target) for each prime condition and model type. We also computed these 95% for human data. We use the overlap in these credible intervals to draw our statistical inferences. If the credible intervals in two priming conditions are non overlapping, we infer that the priming effect in these conditions are different. Similarly, if the credible intervals for predicted priming effect from a model does not overlap with the empirical priming effects, we infer that the model cannot account for human behavior. Such an inference is valid because credible intervals, unlike the frequentist confidence intervals, reflect our confidence about the distribution of the actual effects (so 95% credible interval means that we are 95% sure that the true effect falls within this interval).

### v2

@@ Appendix C Details about the training dataset
Appendix D Exploring model hyperpameters
| Hyperparameter | Equation or Section | Value(s) | Reason |
| Decay () | Eqn 2, Eqn 5 | 0.5 | Vasishth and Engelmann (2021) |
| Latency exponent () | Eqn 3 | 1 | Vasishth and Engelmann (2021) |
| Maximum actiation () | 1.5 | Eqn 4 | Vasishth and Engelmann (2021) |
| Latency factor () | Eqn 3 | Vasishth and Engelmann (2021) | |
| SD of noise distribution () | § 3.2.1 | Vasishth and Engelmann (2021) | |
| Add more noise to retrieve passive. | |||
| # Training sentences | § 5.3 | 0, 100, 1000 | >1000 resulted in almost no passive retrieval. |
| Give up | § 3.2.3 | 100, 1000 | >1000 too much time; 100,1000 same behavior. |
| Reanalysis index () | § 3.2.3 | 1 | Always go back to first word. |
| Entropy weighted sample; SM temp: 1 | Emphasize differences in activation. | ||
| Entropy weighted sample; SM temp: 10 | Make differences in activation more uniform. | ||
| Random seed () | Between 1 to 1280 | Affects training order, random sampling. |
Appendix E Details about statistical models
E.1 Model specification
@@
We fit the maximal model by including all by-participant and by-item random intercepts and slopes. In the case of the predicted data, participant IDs were replaced by model instance IDs.
E.2 Priors
We fit the models using the following weakly informative prior.
@@ E.2 Priors
This prior assumes that the log odds ratio between priming conditions is most likely to be 0 (i.e. no priming effect) and unlikely to be greater than 4 or less than -4. This assumption is based on a meta-analysis of priming in production studies (Mahowald et al., 2016) where the log odds ratio between the prime conditions was not greater than 4 in any of the constructions they considered.
E.3 Statistical inferences for empirical human data
As discussed in the main text, we observed the following qualitative pattern in the proportion of target RRC parses when preceded by different primes: RRC > ProgRRC = FRC > AMV. To ensure that this pattern was statistically valid, we computed Bayes Factors for all of our predictors using the bayestestR package Makowski et al. (2019). We adopt the Bayes Factor scale from Jeffreys (1998) to draw inferences: values greater than 3 and 10 provide moderate and strong evidence for the alternative model, whereas values lower than 0.3 and 0.1 provide moderate and strong evidence for the null model. Therefore, the following Bayes Factor values for our predictors would support the qualitative pattern:
1.
AMV vs. all RCs (C1): > 3
2.
RRC vs. [ProgRRC and FRC] (C2): > 3
3.
ProgRRC vs. FRC (C3): < 0.3
| Predictor | Estimate | 95% CI | Bayes Factor |
| AMV vs. all RCs (C1) | -4.18 | [-5.72, -3.04] | 7.71e+08 |
| RRC vs. [ProgRRC and FRC] (C2) | 0.96 | [0.62, 1.31] | 9.91e+03 |
| ProgRRC vs. FRC (C3) | 0.21 | [-0.18,0.60] | 0.178 |
E.4 Statistical inferences for predicted data
@@
From the posteriors of the Bayesian models, we computed 95% credible intervals for P(RRC prime, target) for each prime condition for the human data, and for each of our model types. If the credible intervals for predicted priming effect from a model do not overlap with the empirical priming effects, we infer that the model cannot account for human behavior. Such an inference is valid because credible intervals, unlike the frequentist confidence intervals, reflect our confidence about the distribution of the actual effects (so 95% credible interval means that we are 95% sure that the true effect falls within this interval).
Appendix F Regular expressions to detect passive responses in the human experiment
We used a three step process to detect passive responses in the human experiment. First, we started with the following regular expression:
[…]
This expression looks for sentences in which the fourth word of the sentence is “by” — all of our target prefixes had only three words (Determiner Noun Verb).
Next we used the following regular expression to detect completions where the fourth word is “by”, but the completion is not passive:
[…]
This expression returns TRUE if the word “by” is followed by just one or two words such as “The thief chased by the dog” or “the thief chased by me”.
Finally, we tagged completions as being passive RRC completion if they matched the first expression and not the second.
Appendix G Limitations
Here we discuss some of the simplifying design decisions we made in SPAWN as a starting point, and their limitations.
Storing discarded categories
As discussed in § 3.2.3, when the parser is regressing to some previous word , it discards all of the categories retrieved from . In the current implementation, SPAWN stores all instances of the discarded categories and uses this to compute inhibition. While storing all instances of the discarded categories is convenient, it is not cognitively plausible. Future work can examine other ways of computing inhibition that relies on summaries of discarded categories, instead of storing all of the instances, and investigate if using summaries results in different priming behavior.
Constraining partial parse states
With our Whiz-Deletion grammar and our current implementation of null element prediction, the model could parse the partial sentence “The cat examined” and end up with an ungrammatical partial parse — i.e., a parse that cannot result in a grammatical continuation — as illustrated below.
| Old Parse state | Word | Retrieved category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/CP | Forward Harmonic Composition | DP/CP |
| DP/CP | NULLwh | CP/(TPDP) | Forward Harmonic Composition | DP/(TPDP) |
| DP/VoiceP | examined | (TPDP)/DP | Forward Harmonic Composition | DP/DP |
This is not a problem in full sentences because this parse state is inconsistent with later words in the sentence, and the model will be forced to re-analyze. However, since our partial target prompts have no additional words, the model could end up with an ungrammatical parse, which is something we assumed would not happen with our human participants. Therefore, we constrained the model such that if it generated a partial state that was not DP/PP or TP/DP, it would be forced to reanalyze. While this is a convenient method to ensure that the model does not end up with an ungrammatical parse, it is unclear if this method accurately models how humans process the partial prompt. Future work can state more explicitly how humans parse the partial sentence such that they are always able to generate grammatical continuations, and then implement this in SPAWN.